# Reliability-Latency-Rate Tradeoff in Low-Latency Communications with Finite-Blocklength Coding

### v3

@@ II-B Effective Bandwidth and Capacity
In this subsection, we will introduce effective bandwidth (EB) and effective capacity (EC). The essence of EB and EC is to use LDT to characterize the tail distribution of the length and delay violation probability. By using the Gartner-Ellis theorem [38] and given a length threshold $L$, QVP satisfies
@@
${\color[rgb]{0,0,0}\alpha_{A}(\theta)=\lim_{M\to+\infty}\frac{1}{\theta M}\ln\mathbb{E}_{A}\left\{e^{\theta A[M]}\right\}.}$ (15)
The EB of the arrival process indicates the minimum constant transmission rate for guaranteeing certain QoS requirements as shown in Eq. (14). Similarly, according to [13], the EC of the service process is defined as
${\color[rgb]{0,0,0}\alpha_{S}(\theta)=\lim_{M\to+\infty}-\frac{1}{\theta M}\ln\mathbb{E}_{S}\left\{e^{-\theta S[M]}\right\}.}$ (16)
The EC of the service process refers to the maximum constant arrival rate that meets the QoS requirements shown in Eq. (14).
@@
Note that the effective rate defined in [14] is equal to $\Lambda(\gamma)$ when $T=1$. In this paper, we use $\Lambda(\gamma)$ to characterize the service-rate gain of the low-latency communication systems.
@@
In Eq. (19), $\eta$ is the probability that the queue is not empty, which can be approximated by the ratio of $\mu_{a}$ to $\mu_{s}$. The notation $a(x)\simeq b(x)$ means that $\lim_{x\to\infty}\frac{a(x)}{b(x)}=1$. In this paper, we focus on the high-load scenario, i.e., $\eta\to 1$, which is more common in emerging applications with the increasing demand for information exchange [39] and is more likely to cause the occurrence of extreme events, e.g., queue length and delay violations  [^5] Note that by setting $\eta=1$ we obtain upper bounds of $\chi$ and $\delta$ defined below with arbitrary loads [40]. Thus, the derived results in this paper can be seen as a converse analysis for the low-load scenario.. $\theta$ is the solution to the following equation [41]:
@@ II-C Main Results
Figure 2: Structures of Section III and Section IV.
In this subsection, we present the main results of this paper. The main results of this paper can be divided into two aspects, which refer to the revised gain-conservation equation with random arrivals in the high-SNR regime and a Laplace’s-method-based EC approximation approach. The detailed analysis and discussion are provided in Sections III and IV. The structure of Sections III and IV are shown in Fig. 2.
Before presenting the main results, we first give some definitions. We define $\zeta=\lim\limits_{\gamma\to+\infty}\frac{\partial{\Lambda(\gamma)}}{\partial{\log_{2}\gamma}}$ as the QoS-constrained service-rate gain, similar to the definition of the service-rate gain in [30]. $\zeta$ indicates the rate of increase of throughput with the increase of $\log_{2}\gamma$ in the high-SNR regime. Let $\varpi=\sqrt{2\log_{2}e}\lim\limits_{\gamma\to+\infty}\frac{\partial\sqrt{-\log_{2}\epsilon(\gamma)}}{\sqrt{N}\partial\log_{2}\gamma}$ denote the reliability gain. $\varpi$ is related to the decay rate of error probability with the increase of $\log_{2}\gamma$ in the high-SNR regime, consistent with the reliability gain in [30]. [^6] Note that the constant $\sqrt{2\log_{2}e}$ is also consistent with [30] since dispersion is treated as a constant in [30]. We can also define $\varpi$ without this constant, which will not affect the main parts of the proposed conclusions. The only difference is that the coefficient of $\varpi$ in the gain-conservation equation will change. We also define $\tau=-\frac{1}{D_{\rm max}}\frac{\partial\log_{2}\delta}{\partial\log_{2}\gamma}$ as the real-time gain, consistent with the definition in [30]. $\tau$ represents the decay rate of DVP with the increase of $\log_{2}\gamma$ in the high-SNR regime. We refer to $\tau$ as real-time gain because real-time usually refers to being controlled and deterministic [3]. $\tau$ represents the decay rate of the probability of extreme events happening. However, $\tau$ can also be named differently, e.g., revised diversity gain  [^7] Diversity gain was originally proposed in [42]. We will present the comparison between $\tau$ and diversity gain in Section III-C., low-latency gain, or timeliness gain. Based on the above definitions, the high-SNR reliability-latency-rate tradeoff is shown in Theorem 1 for low-latency communication systems with FBC and random arrivals. Let $\varrho$ denote an arbitrarily positive constant. The main results and quantities related to gain-conservation equations are summarized in Fig. 3.
@@
Theorem 1. Assume $\epsilon\in(0,0.5]$ is a function of $\gamma$ and $\lim\limits_{\gamma\to+\infty}\epsilon(\gamma)=0$. Under the condition that $\varpi$ is finite, $\theta N=\varrho$, and $N=\Psi(\gamma)$ satisfies
${\color[rgb]{0,0,0}\sup_{\gamma>0}\left|\frac{G\left(\Psi(\gamma),|h[n]|^{2}\gamma,\epsilon(\gamma)\right)}{\Psi(\gamma)}\right|\leq\nu,\quad\text{for all}|h[n]|^{2},}$ (22)
where $\nu\geq 0$ is a finite constant,
  • for a SISO system in an AWGN channel, we have
@@
  • for a SISO system in a Nakagami-$m$ fading channel with a finite $\varpi$, we have
@@
    By combing Eqs. (26) and (27), we have
    $(1-\theta N\log_{2}e)\zeta+c\varpi+\tau=\min\left\{1,\frac{m}{\theta NT\log_{2}e}\right\}.$ (28)
In Eqs. (25) and (28), $c\in[0,1]$ is given by
@@
This theorem summarizes the main results in Section III. The proof of this theorem follows from the proofs of Theorem 5, Theorem 6, Lemma 1, and Theorem 7 (in Appendix D, Appendix E, Appendix F, and Section III-C, respectively).
From the definitions of $\zeta$, $\varpi$, and $\tau$, we find that these three gains characterize the communication system’s performance in terms of the service rate, reliability, and latency with a maximum power constraint. Thus, this theorem presents a fundamental tradeoff among the service rate, reliability, and latency in the high-SNR regime for low-latency communications. We combine Eqs. (26) and (27) to derive (28) so that we can compare Eq. (28) with the Eq. (4) of [30], which has a similar form. In this work, we adopt EC to characterize the service-rate and real-time gains, which is different from the asymptotic analysis in [30]. The main difference between Eqs. (25), (28), and Eq. (4) in [30] results from the influence of random arrivals. Thus, Theorem 1 can be seen as a generalization of Eq. (4) in [30] to the systems with unbounded random arrivals under a large delay threshold. Moreover, [30] lacks a discussion of the assumptions underlying the use of normal approximation in asymptotic analysis, which are provided in this paper.
The assumption shown in Eq. (22) can be met with $G$ function. For simplicity, we omit the index $n$ in the following analysis. Besides, we let $V\left(|h|^{2}\gamma\right)=\left(\log_{2}e\right)^{2}\left(1-\left(1+|h|^{2}\gamma\right)^{-2}\right)$ denote the dispersion term. In [17], Polyanskiy, Poor, and Verdú proposed the upper and lower bounds of $R_{N}^{\epsilon}$, which are given by  [^8] There is a tighter lower bound proposed in [36], which shows that $NC-\sqrt{NV\left(|h|^{2}\gamma\right)}Q^{-1}(\epsilon)+\frac{1}{2}\log_{2}N+O(1)\leq R_{N}^{\epsilon}$. We do not use this bound since the $O(1)$ term in this bound is difficult to analyze. Besides, the lower bound proposed in [18] is sufficient for our analysis.
${\color[rgb]{0,0,0}NC-\sqrt{NV\left(|h|^{2}\gamma\right)}Q^{-1}(\epsilon)+O(1)\leq R_{N}^{\epsilon}\leq NC-\sqrt{NV\left(|h|^{2}\gamma\right)}Q^{-1}(\epsilon)+\frac{\log_{2}N}{2}+O(1),}$
where the upper bound requires $N>\left(\frac{2K\left(|h|^{2}\gamma\right)}{(1-\epsilon)V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}\right)^{2}$ and the lower bound requires $N>\left(\frac{2K\left(|h|^{2}\gamma\right)}{\epsilon V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}\right)^{2}$. Based on the upper and lower bounds of $R_{N}^{\epsilon}$, we can obtain the upper and lower bounds of $G$ function, which are denoted by $g_{u}(|h|^{2}\gamma,\epsilon)$ and $g_{l}(N,|h|^{2}\gamma,\epsilon)$, respectively. Specifically, according to Eq. (632) in [17], the upper bound of $R_{N}^{\epsilon}$ is given by
@@
${\color[rgb]{0,0,0}K\left(|h|^{2}\gamma\right)=c_{0}\left(\frac{|h|^{2}\gamma}{1+|h|^{2}\gamma}\right)^{3}\mathbb{E}_{z}\left\{\left|z^{2}-\frac{2}{\sqrt{|h|^{2}\gamma}}z-1\right|^{3}\right\},}$ (31)
where $c_{0}$ is a positive constant and $z$ is the standard normal variable. $c_{1}$ in Eq. (30) is given by
${\color[rgb]{0,0,0}c_{1}=-2K\left(|h|^{2}\gamma\right)\cdot\min_{y\in\left[1-\epsilon-b_{1},1-\epsilon\right]}\frac{{\rm d}Q^{-1}(y)}{{\rm d}y},}$ (32)
@@
${\color[rgb]{0,0,0}b_{1}=\frac{2K\left(|h|^{2}\gamma\right)}{V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)\sqrt{\left(\frac{2K\left(|h|^{2}\gamma\right)}{V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)(1-\epsilon)}\right)^{2}+1}}.}$ (33)
According to Eqs. (650)-(654) in [17], the lower bound of $R_{N}^{\epsilon}$ is given by
$R_{N}^{\epsilon}$ $\geq NC+\sqrt{NV\left(|h|^{2}\gamma\right)}Q^{-1}(1-\epsilon)+\frac{1}{2}\log_{2}N+{\color[rgb]{0,0,0}g_{l}(N,\gamma,\epsilon)}$ (34)
${\color[rgb]{0,0,0}=\tilde{R}_{N}^{\epsilon}-\frac{1}{2}\log_{2}N+\log_{2}\frac{K\left(|h|^{2}\gamma\right)}{c_{2}V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}-\log_{2}\left[2\left(\frac{\ln 2}{\sqrt{2\pi}}+\frac{2K\left(|h|^{2}\gamma\right)}{V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}\right)\right]}$
${\color[rgb]{0,0,0}\quad\quad\quad\quad\quad\quad+\sqrt{NV\left(|h|^{2}\gamma\right)}\left[Q^{-1}\left(1-\epsilon+\frac{2K\left(|h|^{2}\gamma\right)}{\sqrt{N}V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}\right)+Q^{-1}(\epsilon)\right],}$
where $c_{2}>0$ is a constant. Based on Eqs. (30) and (34), we can obtain the upper and lower bounds of $G(N,|h|^{2}\gamma,\epsilon)$, respectively. Thus, the requirements on $\Psi(\gamma)$ to satisfy Eq. (22) can be determined with the help of Eq. (22) through Eqs. (30) and (34). In Theorems 2 and 3, we present the sufficient conditions with which Eq. (22) holds under different system assumptions.
@@
${\color[rgb]{0,0,0}\Psi(\gamma)>\varsigma_{1},\quad\gamma>0,}$ (35)
where $\Psi(\gamma)$ can be bounded or satisfy $\lim_{\gamma\to+\infty}\Psi(\gamma)=+\infty$. $\varsigma_{1}$ is finite and is given by
${\color[rgb]{0,0,0}\varsigma_{1}=\max_{x\geq 0}\left\{\left(\frac{2K(x)}{(1-\epsilon)V^{\frac{3}{2}}(x)}\right)^{2},\left(\frac{2K(x)}{\epsilon V^{\frac{3}{2}}(x)}\right)^{2}\right\}.}$ (36)
@@
In the proof of Theorem 2, we have also found that, given $N>\varsigma_{1}$ and $\epsilon$, the error term $G(N,\gamma,\epsilon)$ is bounded for $\gamma$, which supplements to [21]. An important future work is to further discuss the case with $N<\varsigma_{1}$.
Theorem 3. With the assumption that $\epsilon\in(0,0.5]$ is a function of $\gamma$ and $\lim\limits_{\gamma\to+\infty}\epsilon(\gamma)=0$, the condition proposed in Eq. (22) is satisfied if $\Psi(\gamma)$ follows
@@
where $\varsigma_{2}$ is a finite and positive constant, satisfying
${\color[rgb]{0,0,0}\varsigma_{2}=4\max_{x\geq 0}\left\{\frac{K^{2}(x)}{V^{3}(x)}\right\}.}$ (38)
@@
See Appendix B.
Theorem 3 presents the specific requirements on $\Psi(\gamma)$ for the cases with $\epsilon(\gamma)\in(0,0.5]$ and $\lim\limits_{\gamma\to+\infty}\epsilon(\gamma)=0$. The form of $\Psi(\gamma)$ is related to $\epsilon(\gamma)$. Given $\epsilon(\gamma)$, $\Psi(\gamma)$ can be conceived according to Eq. (37). For example, let $\Psi(\gamma)=e^{\left[\frac{Q^{-1}(\epsilon(\gamma))}{2}\right]^{2}}+\frac{\varsigma_{2}+1}{\epsilon^{2}(\gamma)}$, which satisfies Eq. (37). Theorem 3 is used to determine the forms of $\Psi(\gamma)$ in Theorems 1, 6, 7 and Corollary 2.
In this paper, we define the constant $\varrho$ as the revised QoS exponent. Different from the definition of $\theta$ in Eq. (14), $\varrho$ is related to the value of the DVP. We can better understand the interpretation of $\varrho$ by providing the expression of the DVP. By combining Eqs. (18), (20), and (21), we obtain
${\color[rgb]{0,0,0}\delta\simeq\exp\left(\frac{D_{\rm max}}{T}\ln\mathbb{E}_{|h|^{2}}\left\{\exp\left(-\varrho T\frac{R_{N}^{\epsilon}[n]}{N}\right)\right\}\right).}$ (39)
From Eq. (39), we find that a larger $\varrho$ results in a smaller $\delta$ given $\gamma$, $\Psi(\cdot)$, and $D_{\rm max}$. Thus, the value of $\varrho$ represents the QoS requirements on DVP.
EC is widely used in research on low-latency communications to evaluate the QoS-constrained throughput of wireless systems. However, it is usually challenging to obtain the closed-form expression of EC, which brings difficulty for the designs of resource allocation and scheduling for maximizing EC. In this paper, we propose a Laplace’s-method-based EC approximation approach, in which normal approximation is used to evaluate the maximum coding rate. Thus, this approximation can be used when normal approximation has satisfying accuracy. For simplicity of expression, let
$r(x)=\log_{2}\left(1+\Xi(x)x\right)-\sqrt{\frac{1}{N}\left(1-(1+\Xi(x)x)^{-2}\right)}Q^{-1}(\epsilon)\log_{2}e+\frac{\log_{2}N}{2N},$ (40)
where $x$ denotes the channel power gain $|h[n]|^{2}$ in time slot $n$. Since $|h[n]|^{2}$ is $i.i.d.$ across time slots, we omit the argument for $x$. $\Xi(x)$ denotes the channel-power-gain-based power allocation scheme. Note that $\Xi(x)$ can also denote the fixed power allocation by setting $\Xi(x)=\gamma$. Besides, let $\ell=\left(\frac{Q^{-1}(\epsilon)}{\sqrt{N}}\right)^{2}$.
Theorem 4. For a low-latency communication system that adopts a power allocation scheme $\Xi(x)$ satisfying $\Xi(x)+\frac{{\rm d}\Xi(x)}{{\rm d}x}x\geq 0$ with channel distribution $f(\cdot)$, EC is approximated as
$\alpha_{S}(\theta)=Nr(x^{*})-\frac{\ln f(x^{*})+\frac{1}{2}\ln 2\pi-\frac{1}{2}\ln|r^{\prime\prime}(x^{*})|-\frac{1}{2}\ln\theta NT}{\theta T}+O\left(\frac{1}{\theta^{2}T^{2}N}\right),$ (41)
where $x^{*}$ is given by
@@
Theorem 4 summarizes the main results in Section IV. The proof of this theorem follows from the proof of Lemma 2 and Theorem 8 (in Appendix H and I, respectively).
The approximation proposed in Theorem 4 gives us a way to obtain the closed-form expression of the EC, which can help us further analyze the QoS-constrained performance of low-latency communication systems and compare different scheduling schemes. The condition mentioned in Theorem 4 can be satisfied by some classical allocation schemes. More detailed discussion on $\Xi(x)$ is presented in Section IV-B.

### v4

@@ II-B Effective Bandwidth and Capacity
In this subsection, we will introduce effective bandwidth (EB) and effective capacity (EC). The essence of EB and EC is to use large deviation theory (LDT) to characterize the tail distribution of the length and delay violation probability. By using the Gartner-Ellis theorem [38] and given a length threshold $L$, QVP satisfies
@@
$\alpha_{A}(\theta)=\lim_{M\to+\infty}\frac{1}{\theta M}\ln\mathbb{E}_{A}\left\{e^{\theta A[M]}\right\}.$ (15)
EB of the arrival process indicates the minimum constant transmission rate for guaranteeing certain QoS requirements as shown in Eq. (14). Similarly, according to [13], the EC of the service process is defined as
$\alpha_{S}(\theta)=\lim_{M\to+\infty}-\frac{1}{\theta M}\ln\mathbb{E}_{S}\left\{e^{-\theta S[M]}\right\}.$ (16)
EC of the service process refers to the maximum constant arrival rate that meets the QoS requirements shown in Eq. (14).
@@
$\Lambda(\gamma)$ $=-\frac{1}{\theta NT}\ln\mathbb{E}_{|h|^{2}}\left\{e^{-\theta Ts[n]}\right\}$ (18)
@@
Note that the effective rate defined in [14] is equal to $\Lambda(\gamma)$ when $T=1$. In this paper, we use $\Lambda(\gamma)$ to characterize the service-rate gain of the low-latency communication systems, which aligns with [14].
@@
In Eq. (19), $\eta$ is the probability that the queue is not empty, which can be approximated by the ratio of $\mu_{a}$ to $\mu_{s}$. [^5] Note that by setting $\eta=1$ we obtain upper bounds of $\chi$ and $\delta$ defined below with arbitrary loads [39] Thus, the derived results in this paper can be seen as a converse analysis for the low-load scenario. The notation $a(x)\simeq b(x)$ means that $\lim_{x\to\infty}\frac{a(x)}{b(x)}=1$. In this paper, we focus on the high-load scenario, i.e., $\eta\to 1$, which is more common in emerging applications with the increasing demand for information exchange [40] and is more likely to cause the occurrence of extreme events, e.g., queue length and delay violations. $\theta$ is the solution to the following equation [41]:
@@ II-C Main Results
In this subsection, we present the main results of this paper. The main results of this paper can be divided into two aspects, which refer to the revised gain-conservation equations with random arrivals in the high-SNR regime and a Laplace’s-method-based approximation approach for EC. Detailed analysis and discussions are provided in Sections III and IV. The structures of Sections III and IV are illustrated in Fig. 2.
Before presenting the main results, we first introduce key definitions. We define $\zeta=\lim\limits_{\gamma\to+\infty}\frac{\partial{\Lambda(\gamma)}}{\partial{\log_{2}\gamma}}$ as the QoS-constrained service-rate gain, similar to the concept of the service-rate gain in [30]. $\zeta$ indicates the growth rate of throughput with respect to $\log_{2}\gamma$ in the high-SNR regime. Next, we define the reliability gain $\varpi=\sqrt{2\log_{2}e}\lim\limits_{\gamma\to+\infty}\frac{\partial\sqrt{-\log_{2}\epsilon(\gamma)}}{\sqrt{N}\partial\log_{2}\gamma}$, which quantifies the decay rate of error probability as $\log_{2}\gamma$ increases in the high-SNR regime, consistent with the reliability gain proposed in [30]. [^6] Note that the constant $\sqrt{2\log_{2}e}$ aligns with [30] since dispersion term is treated as a constant in [30]. Omitting this constant does not alter the conclusions but would modify the coefficient of $\varpi$ in the gain-conservation equations. We also define the real-time gain $\tau=-\frac{1}{D_{\rm max}}\frac{\partial\log_{2}\delta}{\partial\log_{2}\gamma}$, which is analogous to the definition in [30]. $\tau$ captures the decay rate of DVP with increasing $\log_{2}\gamma$ in the high-SNR regime. We call $\tau$ the real-time gain since real-time usually refers to being controlled and deterministic [3]. $\tau$ represents the decay rate of the probability of extreme events happening. Alternatively, $\tau$ can also be named differently, e.g., revised diversity gain  [^7] Diversity gain was initially proposed in [42]. A detailed comparison between $\tau$ and diversity gain is provided in Section III-C., low-latency gain, or timeliness gain.
Based on the above definitions of $\zeta$, $\varpi$, and $\tau$, Theorem 1 presents the high-SNR reliability-latency-rate tradeoff through revised gain-conservation equations for low-latency systems with FBC and random arrivals. For simplicity, let $\varrho$ denote an arbitrarily positive constant, which is named as revised QoS exponent. The main results and terms related to the gain-conservation equations are illustrated in Fig. 3.
@@
Theorem 1. Let $\epsilon\in(0,0.5]$ be a function of $\gamma$ with $\lim\limits_{\gamma\to+\infty}\epsilon(\gamma)=0$. Additionally, assume that $\varpi$ is finite, $\theta N=\varrho$, and $N=\Psi(\gamma)$, where $\Psi(\gamma)$ satisfies the following condition:
$\sup_{\gamma>0}\left|\frac{G\left(\Psi(\gamma),|h[n]|^{2}\gamma,\epsilon(\gamma)\right)}{\Psi(\gamma)}\right|\leq\nu,\quad\text{for all }|h[n]|^{2}.$ (22)
In Eq. (22), $\nu\geq 0$ is a finite constant.
  • For a SISO system over an AWGN channel, the following relations hold:
@@
  • For a SISO system over a Nakagami-$m$ fading channel, the following relations hold:
@@
    By combining Eqs. (26) and (27), we obtain
    $(1-\theta N\log_{2}e)\zeta$ $+c\varpi+\tau$ (28)
    $=\min\left\{1,\frac{m}{\theta NT\log_{2}e}\right\}.$
In Eqs. (23), (25), (26) and (28), $c\in[0,1]$ is given by
@@
This theorem summarizes the main results in Section III. The proof of this theorem is derived from the proofs of Theorem 5, Theorem 6, Lemma 1, and Theorem 7, as detailed in Appendix D, Appendix E, Appendix F, and Section III-C, respectively. ∎
@@
From the definitions of $\zeta$, $\varpi$, and $\tau$, we find that these three gains effectively characterize the performance of communication systems in terms of the service rate, reliability, and latency under a maximum power constraint. Consequently, this theorem presents a fundamental tradeoff among the service rate, reliability, and latency in the high-SNR regime for low-latency communications. By combining Eqs. (26) and (27), we derive (28), which allows for a comparison with Eq. (4) from [30], as they exhibit a similar form. In this work, we adopt EC to characterize the service-rate and real-time gains, which is different from the method utilized for the asymptotic analysis in [30]. The primary difference between Eqs. (25), (28), and Eq. (4) in [30] results from the influence of random arrivals. Therefore, Theorem 1 can be regarded as a generalization of Eq. (4) in [30] for systems with unbounded random arrivals under a large delay threshold. Furthermore, [30] lacks a discussion of the assumptions underlying the use of normal approximation in asymptotic analysis, which are provided in this paper.
The assumption presented in Eq. (22) can be met with certain $G$ functions. We provide the discussion on the requirements for $G$ function in the following. For simplicity, we omit the index $n$ in the subsequent analysis. Additionally, we let $V\left(|h|^{2}\gamma\right)=\left(\log_{2}e\right)^{2}\left(1-\left(1+|h|^{2}\gamma\right)^{-2}\right)$ denote the dispersion term, which is a function of $|h|^{2}\gamma$. In [17], Polyanskiy, Poor, and Verdú proposed the upper and lower bounds for $R_{N}^{\epsilon}$, which are given by  [^8] There is a tighter lower bound proposed in [36], which shows that $NC-\sqrt{NV\left(|h|^{2}\gamma\right)}Q^{-1}(\epsilon)+\frac{1}{2}\log_{2}N+O(1)\leq R_{N}^{\epsilon}$. However, we do not use this bound due to the complexity of analyzing the $O(1)$ term in this bound. Moreover, the lower bound proposed in [18] is sufficient for our analysis.
$NC$ $-\sqrt{NV\left(|h|^{2}\gamma\right)}Q^{-1}(\epsilon)+O(1)\leq R_{N}^{\epsilon}$ (30)
$\leq NC-\sqrt{NV\left(|h|^{2}\gamma\right)}Q^{-1}(\epsilon)+\frac{\log_{2}N}{2}+O(1),$
where the upper bound requires
$N>\left(\frac{2K\left(|h|^{2}\gamma\right)}{(1-\epsilon)V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}\right)^{2},$
and the lower bound requires
$N>\left(\frac{2K\left(|h|^{2}\gamma\right)}{\epsilon V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}\right)^{2}.$
Based on the upper and lower bounds of $R_{N}^{\epsilon}$, we can derive the upper and lower bounds of $G$ function, which are denoted by $g_{u}(|h|^{2}\gamma,\epsilon)$ and $g_{l}(N,|h|^{2}\gamma,\epsilon)$, respectively. Specifically, according to Eq. (632) in [17], the analytical form of the upper bound of $R_{N}^{\epsilon}$ is given in Eq. (31), which is presented at the bottom of the last page.
@@
$K\left(|h|^{2}\gamma\right)=$ $c_{0}\left(\frac{|h|^{2}\gamma}{1+|h|^{2}\gamma}\right)^{3}\cdot$ (32)
$\quad\quad\quad\mathbb{E}_{z}\left\{\left|z^{2}-\frac{2}{\sqrt{|h|^{2}\gamma}}z-1\right|^{3}\right\},$
where $c_{0}$ is a positive constant and $z$ represents a standard normal variable.
$c_{1}$ in Eq. (31) is given by
$c_{1}=-2K\left(|h|^{2}\gamma\right)\cdot\min_{y\in\left[1-\epsilon-b_{1},1-\epsilon\right]}\frac{{\rm d}Q^{-1}(y)}{{\rm d}y},$ (33)
@@
$b_{1}=\frac{2K\left(|h|^{2}\gamma\right)}{V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)\sqrt{\left(\frac{2K\left(|h|^{2}\gamma\right)}{V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)(1-\epsilon)}\right)^{2}+1}}.$ (34)
According to Eqs. (650)-(654) in [17], the lower bound of $R_{N}^{\epsilon}$ is expressed in Eq. (35) (located at the bottom of this page), where $c_{2}>0$ is a constant. Based on Eqs. (31) and (35), we can derive the upper and lower bounds of $G(N,|h|^{2}\gamma,\epsilon)$, respectively. Consequently, the requirements for $\Psi(\gamma)$ to satisfy Eq. (22) can be determined using Eqs. (31) and (35). In Theorems 2 and 3, we present the sufficient conditions under which Eq. (22) holds for different system assumptions.
$R_{N}^{\epsilon}$ $\geq NC+\sqrt{NV\left(|h|^{2}\gamma\right)}Q^{-1}(1-\epsilon)+\frac{1}{2}\log_{2}N+g_{l}(N,\gamma,\epsilon)$ (35)
$=\tilde{R}_{N}^{\epsilon}-\frac{1}{2}\log_{2}N+\log_{2}\frac{K\left(|h|^{2}\gamma\right)}{c_{2}V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}-\log_{2}\left[2\left(\frac{\ln 2}{\sqrt{2\pi}}+\frac{2K\left(|h|^{2}\gamma\right)}{V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}\right)\right]$
$\quad\quad\quad\quad\quad\quad+\sqrt{NV\left(|h|^{2}\gamma\right)}\left[Q^{-1}\left(1-\epsilon+\frac{2K\left(|h|^{2}\gamma\right)}{\sqrt{N}V^{\frac{3}{2}}\left(|h|^{2}\gamma\right)}\right)+Q^{-1}(\epsilon)\right].$
@@
$\Psi(\gamma)>\varsigma_{1},\quad\text{for all }\gamma>0,$ (36)
where $\Psi(\gamma)$ may either be bounded or satisfy $\lim_{\gamma\to+\infty}\Psi(\gamma)=+\infty$. $\varsigma_{1}$ is finite and is given by
@@
Furthermore, the proof of Theorem 2 reveals that, for $N>\varsigma_{1}$ and a given $\epsilon$, the error term $G(N,\gamma,\epsilon)$ remains bounded with respect to $\gamma$, which provides additional insights complementary to those in [21]. An important future work is to further investigate the case where $N<\varsigma_{1}$.
Theorem 3. Under the assumption that $\epsilon\in(0,0.5]$ is a function of $\gamma$ and $\lim\limits_{\gamma\to+\infty}\epsilon(\gamma)=0$, the condition proposed in Eq. (22) is satisfied if $\Psi(\gamma)$ follows
@@
where $\varsigma_{2}$ is a finite and positive constant defined as
$\varsigma_{2}=4\max_{x\geq 0}\left\{\frac{K^{2}(x)}{V^{3}(x)}\right\}.$ (39)
@@
See Appendix B. ∎
Theorem 3 outlines the specific requirements for $\Psi(\gamma)$ under the conditions that $\epsilon(\gamma)\in(0,0.5]$ and $\lim\limits_{\gamma\to+\infty}\epsilon(\gamma)=0$. The form of $\Psi(\gamma)$ is related to $\epsilon(\gamma)$. Given a specific form of $\epsilon(\gamma)$, $\Psi(\gamma)$ can be formulated according to Eq. (38). For example, one can let $\Psi(\gamma)=e^{\left[\frac{Q^{-1}(\epsilon(\gamma))}{2}\right]^{2}}+\frac{\varsigma_{2}+1}{\epsilon^{2}(\gamma)}$, which satisfies Eq. (38). Theorem 3 serves as a foundation for determining the forms of $\Psi(\gamma)$ utilized in Theorems 1, 6, 7 and Corollary 2.
In this paper, we define the constant $\varrho$ as the revised QoS exponent. Different from the definition of $\theta$ in Eq. (14), $\varrho$ is associated with the value of the DVP. We can gain a clearer understanding of the interpretation of $\varrho$ by presenting the expression of the DVP. By combining Eqs. (18), (20), and (21), we obtain
$\delta\simeq\exp\left(\frac{D_{\rm max}}{T}\ln\mathbb{E}_{|h|^{2}}\left\{\exp\left(-\varrho T\frac{R_{N}^{\epsilon}[n]}{N}\right)\right\}\right).$ (40)
From Eq. (40), we observe that a larger $\varrho$ results in a smaller $\delta$ given $\gamma$, $\Psi(\cdot)$, and $D_{\rm max}$. Therefore, the value of $\varrho$ can represent the QoS requirements for the DVP.
EC is widely utilized in research on low-latency communications to assess the QoS-constrained throughput of wireless systems. However, it is usually challenging to obtain a closed-form expression of EC, which brings difficulty for the designs of resource allocation and scheduling strategies aimed at maximizing EC. In this paper, we propose a Laplace’s-method-based EC approximation approach, in which normal approximation is employed to evaluate the maximum coding rate. Thus, this approximation is applicable when the normal approximation has satisfying accuracy. For the simplicity of expression, let
$r(x)=$ $\log_{2}\left(1+\Xi(x)x\right)-$ (41)
$\sqrt{\frac{1}{N}\left(1-\frac{1}{(1+\Xi(x)x)^{2}}\right)}Q^{-1}(\epsilon)\log_{2}e+\frac{\log_{2}N}{2N},$
where $x$ denotes the channel power gain $|h[n]|^{2}$ in time slot $n$. Since $|h[n]|^{2}$ is $i.i.d.$ across time slots, we omit the argument for $x$. The function $\Xi(x)$ denotes the channel-power-gain-based power allocation scheme. Note that $\Xi(x)$ can also denote the fixed power allocation by setting $\Xi(x)=\gamma$. Besides, let $\ell=\left(\frac{Q^{-1}(\epsilon)}{\sqrt{N}}\right)^{2}$.
Theorem 4. For a low-latency communication system that adopts a power allocation scheme $\Xi(x)$ satisfying $\Xi(x)+\frac{{\rm d}\Xi(x)}{{\rm d}x}x\geq 0$ with the channel distribution $f(\cdot)$, EC is approximated using Eq. (42), which is presented at the bottom of this page. In Eq. (42), $x^{*}$ is given by
@@
$\alpha_{S}(\theta)=Nr(x^{*})-\frac{\ln f(x^{*})+\frac{1}{2}\ln 2\pi-\frac{1}{2}\ln|r^{\prime\prime}(x^{*})|-\frac{1}{2}\ln\theta NT}{\theta T}+O\left(\frac{1}{\theta^{2}T^{2}N}\right).$ (42)
@@
Theorem 4 summarizes the main results in Section IV. The proof of this theorem follows from the proofs of Lemma 2 and Theorem 8 (in Appendix H and I, respectively). ∎
The approximation proposed in Theorem 4 provides a method to obtain a closed-form expression for EC, which facilitates further analysis of the QoS-constrained performance of low-latency communication systems and allows for theoretical comparison of different scheduling schemes. The condition stated in Theorem 4 can be satisfied by various classical allocation schemes. More detailed discussions on $\Xi(x)$ are presented in Section IV-B.